OpenStudy (anonymous):

how do you solve \sqrt{3k-11}=\sqrt{5-k}

OpenStudy (anonymous):

doe sit look like this: \[\sqrt{3k-11}=\sqrt{5-k}\]

OpenStudy (anonymous):

yes im sorry cant do that on my computer

OpenStudy (anonymous):

it's ok k=4 then:)

OpenStudy (anonymous):

oh i got 16 because i dont know how to do the steps

OpenStudy (anonymous):

would you help me write the steps out

OpenStudy (anonymous):

To be honest I used Mathway problem solver... I can't remember how to solve each step exactly.... let me think though

OpenStudy (anonymous):

ok if you cant its ok thank you

OpenStudy (anonymous):

Ya idk do you have answer choices?

OpenStudy (anonymous):

i had got 16 because i dont get it

OpenStudy (anonymous):

oh but it's not a multiple answer choice?

OpenStudy (anonymous):

no its solving equation worksheet its square root equation

OpenStudy (anonymous):

ok... well you could solve by factoring

OpenStudy (anonymous):

i was thinking that

OpenStudy (anonymous):

ya you can solve it by factoring